OpenStudy (anonymous):

here we go again lol: 5/x + 4/x-3 =1 A quadratic! please please please help me, thanks so much!!! :)

OpenStudy (anonymous):

is this \[\frac{5}{x}+\frac{4}{x-3}=1\]?

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

\[5(x-3)+4x=x(x-3)\] \[x^2-12 x+15 = 0\] then the quadratic formula

OpenStudy (anonymous):

do you know the formula? i cant find my paper :(

OpenStudy (anonymous):

you can find it on line in many places, for example http://www.sosmath.com/algebra/quadraticeq/quadraformula/quadraformula.html

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

do you know how to use it?

OpenStudy (anonymous):

can someone factor this for me, show me how to do it that way, please

OpenStudy (anonymous):

instead of the quadratic formula?

OpenStudy (anonymous):

it does not factor. you have to use \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

that is it does not factor using integers.

OpenStudy (anonymous):

satellite: wait a second for her to reply

OpenStudy (anonymous):

you can complete the square as follows \[x^2-12x=-15\] \[(x-6)^2=-15+36=21\] \[x-6=\pm\sqrt{21}\] \[x=6\pm\sqrt{21}\]

OpenStudy (anonymous):

ok just a moment....

OpenStudy (anonymous):

that looks right, but it wants another # with the 6

OpenStudy (anonymous):

nevermind, it took it, thanks!!!!! :D :D